OpenStudy (anonymous):

What is the slope-intercept form of the function described by this table? x 1 2 3 4 y 2 7 12 17 y = x +

sammixboo (sammixboo):

Doyou know the formula for slope-intercept form?

sammixboo (sammixboo):

y = mx + b is the formula. m = slope y = y value x =x value b = y intercept (starting point) Well b is given.It is 2, so now we have y = mx + 2 Now, let's find the slope. Do you know the formula tofind the slope given two points?

OpenStudy (anonymous):

no, i'm not good with function-slopes

sammixboo (sammixboo):

Ah, OK. Well the formula is \(\dfrac{y_2 - y_1}{x_2 - x_1}\) Let's use two of the points (1, 2) and (2, 7) (1, 2) 1 = \(x_1\) 2 = \(y_1\) (2, 7) 2 = \(x_2\) 7 = \(y_2\) Plug in the values.. \(slope = \dfrac{2-1}{7-2}\)

sammixboo (sammixboo):

What is 2 -1 and 7 - 2?

OpenStudy (anonymous):

1,5

OpenStudy (anonymous):

@sammixboo